PLEASE HELP ASAP!
Find the amount accumulated FV in the sinking fund. Round your answer to the nearest cent.
$350 is deposited monthly for 15 years at 6% per year
FV= ?

Answer by Theo(13342)

if you deposit the money at the end of the month, then you need this formula.

FUTURE VALUE OF AN ANNUITY WITH END OF TIME PERIOD PAYMENTS

if you deposit the money at the beginning of the month, then you need this formula.

FUTURE VALUE OF AN ANNUITY WITH BEGINNING OF TIME PERIOD PAYMENTS

both these formulas are found in the following link.

https://www.algebra.com/algebra/homework/Finance/THEO-2016-04-29.lesson#formulas

you can do it manually or you can use the following onlikne calcujlator that will do it for you.

http://arachnoid.com/finance/index.html

i used the online calculator.

i did end of time period first and then beginning of time period.

following are the results.

first is end of month payments.
fv = 101786.55
$$$

second is beginning of month payments.
fv = 102295.48
$$$

you enter payment of 350.
you enter pv = 0 and fv = 0
you enter number of time periods = 15 * 12 = 180.
you enter interest rate percent = 6/12 = .5
you select payment at beginning or end of time period.
you click on fv.
you get your answer.